# Data-Driven Learnability Transition of Measurement-Induced Entanglement

### v1

@@ Abstract
Motivated by the recent success of data-driven quantum learning methods based solely on measurement records [54, 55, 56], we ask in this Letter: under what conditions can MIE be learned with polynomial resources from data alone—i.e., without post-selection and without any prior knowledge of state preparation? To address this question, we study the entanglement generated between two distant qubits $A$ and $B$ after all other qubits in a many-body state are measured. Using measurement outcomes only, we train a transformer-based neural network in a self-supervised manner to estimate the post-measurement state on $AB\equiv A\cup B$ conditioned on a given measurement outcome. Instead of reconstructing MIE directly, we estimate the total entanglement entropy of $AB$, which we use as an uncertainty metric quantifying the learnability of MIE. We focus on two settings—random one-dimensional (1D) all-to-all circuits and two-dimensional (2D) nearest-neighbor circuits—and in both we observe a clear learnability transition with increasing circuit depth. In the learnable phase, the uncertainty decreases with additional measurement data and model parameters, reflecting increasing time and space complexity. In the unlearnable phase, the uncertainty saturates or even grows. We further probe this learnability transition in the presence of noise through simulations and experiments on IBM QPU ibm_marrakesh, finding that the transition remains observable under realistic noise. Overall, our results demonstrate the power of data-driven methods to extract informative features of the post-measurement state that reveal MIE, while simultaneously clarifying the fundamental limitations of learning MIE from measurement data alone.
@@
which utilizes $S_{AB,m}\geq 0$. Thus, the true entanglement $S_{A,m}$ lies within an interval of width $S_{AB,m}^{\text{QC}}$ for each outcome $m$. Nevertheless, direct computing $S_{AB,m}^{\text{QC}}$ requires complete knowledge of $\sigma_{AB,m}$. To access this quantity, we employ classical shadow tomography [57]. Qubits $A$ and $B$ are measured in random single-qubit Pauli basis, producing classical snapshots $\sigma_{AB,m}^{s}$ that is an unbiased estimator satisfying $\mathbb{E}_{s}[\sigma_{AB,m}^{s}]=\sigma_{AB,m}$, where $\mathbb{E}_{s}$ denotes averaging over different snapshots. We then define
@@
This learnability transition directly reflects an intrinsic computational phase transition in random quantum circuits, closely related to a finite-depth many-body teleportation transition [61]. As shown in the inset of Fig. 1(d), the MIE between $A$ and $B$ exhibits a finite-depth transition. Below a critical depth, the dynamics remain in a non-teleporting regime, correlations are essentially local, MIE vanishes as the system size becomes larger, and efficient classical algorithms exist for approximate random-circuit sampling and the dynamics exhibit low effective complexity [62]. In this regime, states are expected to be PAC-learnable [63], implying that suitably designed neural networks can achieve accurate estimation with polynomial data and parameters resources [64]—the learnable phase. Beyond the critical depth, the circuit enters a teleporting regime where long-range entanglement emerges. Efficient classical simulation is believed to fail here, rendering average-case random-circuit sampling computationally intractable and establishing the basis of quantum computational advantage [65]. Therefore, polynomial-resource machine-learning models are not expected to succeed—defining the unlearnable phase. From this perspective, the learnability transition provides a direct machine-learning probe of computational complexity, faithfully tracking the underlying computational phase boundary. To demonstrate the finite-depth nature of this transition, we compare the results across different system sizes $L=24$ using the same $N_{p}$ and $N_{m}$. At small $t$, $\Delta$ for $L=24$ is comparable to or smaller than that for $L=20$. As $t$ increases, the curves cross, indicating that the transition occurs at a finite depth. Accurately determining the critical depth would require simulations at larger $L$ to minimize finite-size effects.
@@
Discussions—Our results have several implications for extracting MIE in practice and for understanding fundamental limits of quantum learning. In the shallow-depth regime, the observed learnability indicates that neural networks trained solely on measurement data can serve as surrogates for classical simulation whenever the latter remains tractable. This demonstrates the possibility of deploying data-driven neural network decoders instead of simulation-based decoders for quantum error correction [71, 72, 73, 74, 75]. By contrast, the loss of learnability at large depth provides evidence for quantum advantage: when classical simulation becomes intractable, machine-learning approach likewise fail to recover MIE with polynomial resources. This further suggests that in deep circuits within our setup, not only is substantial MIE generated, but crucially, this entanglement remains undetectable. Harnessing MIE in such regimes—rather than merely witnessing its presence—thus remains an important outstanding challenge.
Our work also suggests several interesting directions for future investigation. The present resource analysis accounts for quantum time complexity via $N_{m}$ and classical space complexity through $N_{p}$, but does not incorporate the classical classical time complexity of training itself. Including training overhead would sharpen the boundary of efficient MIE extraction and is essential for a complete complexity-theoretic characterization. In addition, our study has focused on states generated by random circuits, which represent an average-case scenario. Recent work [76] has explored data-driven learning of MIE in structured states such as GHZ and cluster states—where a similar learnability phase transition arises in the latter when tuning the measurement direction. Since physically motivated states—e.g., ground states of interacting Hamiltonians or states with specific symmetries—typically exhibit structure absent in random ensembles, exploring learnability transitions in such settings is a natural and important next step. We leave this to future work.
@@ Acknowledgements.
Acknowledgments—We acknowledge Yi-Zhuang You for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission.

### v2

@@ Abstract
Motivated by the recent success of data-driven quantum learning methods based solely on measurement records [54, 55, 56], we ask in this Letter: under what conditions can MIE be learned with polynomial resources from data alone—i.e., without post-selection and without any prior knowledge of state preparation? To address this question, we study the entanglement generated between two distant qubits $A$ and $B$ after all other qubits in a many-body state are measured. Using measurement outcomes only, we train a transformer-based neural network in a self-supervised manner to estimate the post-measurement state on $AB\equiv A\cup B$ conditioned on a given measurement outcome. Instead of reconstructing MIE directly, we estimate the total entanglement entropy of $AB$, which we use as an uncertainty metric quantifying the learnability of MIE. We focus on two settings—random one-dimensional (1D) all-to-all circuits and two-dimensional (2D) nearest-neighbor circuits—and in both we observe a clear learnability transition with increasing circuit depth. In the learnable phase, the uncertainty decreases with additional measurement data and model parameters, reflecting increasing time and space complexity. In the unlearnable phase, the uncertainty saturates or even grows. We further probe this learnability transition in the presence of noise through simulations and experiments on IBM QPU ibm_marrakesh, finding that the transition remains observable under realistic noise. Overall, our results demonstrate the power of data-driven methods to extract informative features of the post-measurement state that reveal MIE, while simultaneously clarifying the fundamental limitations of learning MIE from measurement data alone. We note that a recent study [57] investigated MIE learnability for GHZ and cluster states using a similar data-driven approach, where signatures of a learnability transition were observed for the latter upon tuning the measurement direction. In contrast, our work addresses generic states generated by random quantum circuits and provides an explicit characterization of the emergence of distinct learnable and unlearnable regimes.
@@
which utilizes $S_{AB,m}\geq 0$. Thus, the true entanglement $S_{A,m}$ lies within an interval of width $S_{AB,m}^{\text{QC}}$ for each outcome $m$. Nevertheless, directly computing $S_{AB,m}^{\text{QC}}$ requires complete knowledge of $\sigma_{AB,m}$. To access this quantity, we employ classical shadow tomography [58]. Qubits $A$ and $B$ are measured in random single-qubit Pauli basis, producing classical snapshots $\sigma_{AB,m}^{s}$ that are an unbiased estimators satisfying $\mathbb{E}_{s}[\sigma_{AB,m}^{s}]=\sigma_{AB,m}$, where $\mathbb{E}_{s}$ denotes averaging over different snapshots. We then define
@@
This learnability transition directly reflects an intrinsic computational phase transition in random quantum circuits, closely related to a finite-depth many-body teleportation transition [62]. As shown in the inset of Fig. 1(d), the MIE between $A$ and $B$ exhibits a finite-depth transition. Below a critical depth, the dynamics remain in a non-teleporting regime, correlations are essentially local, MIE vanishes as the system size becomes larger, and efficient classical algorithms exist for approximate random-circuit sampling and the dynamics exhibit low effective complexity [63]. In this regime, states are expected to admit efficient learning [64], implying that suitably designed neural networks can achieve accurate estimation with polynomial data and parameters resources [65]—the learnable phase. Beyond the critical depth, the circuit enters a teleporting regime where long-range entanglement emerges. Efficient classical simulation is believed to fail here, rendering average-case random-circuit sampling computationally intractable and establishing the basis of quantum computational advantage [66]. Therefore, polynomial-resource machine-learning models are not expected to succeed—defining the unlearnable phase. From this perspective, the learnability transition provides a direct machine-learning probe of computational complexity, faithfully tracking the underlying computational phase boundary. To demonstrate the finite-depth nature of this transition, we compare the results across different system sizes $L=24$ using the same $N_{p}$ and $N_{m}$. At small $t$, $\Delta$ for $L=24$ is comparable to or smaller than that for $L=20$. As $t$ increases, the curves cross, indicating that the transition occurs at a finite depth. Accurately determining the critical depth would require simulations at larger $L$ to minimize finite-size effects.
@@
Discussions—Our results have several implications for extracting MIE in practice and for understanding fundamental limits of quantum learning. In the shallow-depth regime, the observed learnability indicates that neural networks trained solely on measurement data can serve as surrogates for classical simulation whenever the latter remains tractable. This demonstrates the possibility of deploying data-driven neural network decoders instead of simulation-based decoders for quantum error correction [72, 73, 74, 75, 76]. By contrast, the loss of learnability at large depth provides evidence for quantum advantage: when classical simulation becomes intractable, machine-learning approaches likewise fail to recover MIE with polynomial resources. This further suggests that in deep circuits within our setup, not only is substantial MIE generated, but crucially, this entanglement remains undetectable. Harnessing MIE in such regimes—rather than merely witnessing its presence—thus remains an important outstanding challenge.
Our work also suggests several interesting directions for future investigation. The present resource analysis accounts for quantum time complexity via $N_{m}$ and classical space complexity through $N_{p}$, but does not incorporate the classical time complexity of training itself. Including training overhead would sharpen the boundary of efficient MIE extraction and is essential for a complete complexity-theoretic characterization. In addition, our study has focused on states generated by random circuits, which represent an average-case scenario. Since physically motivated states—e.g., ground states of interacting Hamiltonians or states with specific symmetries—typically exhibit structure absent in random ensembles, exploring learnability transitions in such settings is a natural and important next step. We leave this to future work.
@@ Acknowledgements.
Acknowledgments—We acknowledge Yi-Zhuang You and Pengfei Zhang for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission.